4. What is the pH of a buffer solution that is prepared by making a solution 0.25 M NH, and 0.15 M NH?
Answer #1

[NH3] = 0.25M

[NH4^+]. = 0.15M

PKb. = 4.75

POH. =. PKb +log[NH4^+]/[NH3]

= 4.75 +log0.15/0.25

= 4.75 -0.2218

= 4.5282

PH. = 14-POH

=9.4718